Let R be the set of real numbers and the mapping f :R ? R and g : R ? R be defined by f(x) = 5 – x2
and g(x) = 3x – 4 , then
the value of (fog)(–1) is
(A) –44 (B) –54 (C) –32 (D) –64

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: -44
f(x) = 5-x^{2}
g(x) = 3x-4
fo(g(x)) = 5-(3x-4)^{2}
fo(g(-1)) = 5-(3(-1)-4)^{2} = 5 - (-7)^{2} = 5-49 = -44
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty